Sonics and gas analyzers

v07 4m

The plan is to replace it with the last RMYoung. 2% bad samples all the time would be better than 2% good samples during the day.

tse01 10m

CSAT3A on port 2 is reporting all bad samples as of 2017-06-02,08 UTC. So we only have to replace whichever of head or box is bad, or both, but we have spares. Cycling power did not help.

tnw09 10m

Lots of flagged samples still. This is a CSAT3 IRGA. Up to %30 of the winds are flagged in qctables, need to check on the gas diagnostic.

tnw07b 4m

Diagnostic still mostly bad.

TRH Sensors

tse11 60m

Still needs to be replaced.

Other issues

tse05 power

Isabel and I installed a second solar panel at tse05 around 16z. The panel is more horizontal and is aimed more west, since the first panel is aimed more east. It Looks like voltage jumped up a tenth, but maybe there will not be enough charging left today to get it through tonight. Maybe tomorrow night.

  • No labels

1 Comment

  1. At tse05, at least P and Vdsm have offscale values during the nights of May 31, Jun 1,2,3 presumably due to the low voltage.  We'll have to remove these values in post-processing (probably just by a maxVale parameter in the .xml).